# Measurement Error and Counterfactuals in Quantitative Trade and Spatial Models

### v4

@@ 5 Applications
In this section I discuss the applications in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). In both cases, accounting for estimation and measurement error leads to substantial uncertainty around the counterfactual predictions.
5.1 Application 1: Adao, Costinot, and Donaldson (2017)
@@ 5.1.1 Model and Counterfactual Question of Interest
The empirical application of Adao, Costinot, and Donaldson (2017) investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs had stayed constant at their 1995 levels. They consider $n$ countries and $T$ time periods.
The counterfactual objects of interest is the change in China’s welfare, defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change where China’s trade costs are fixed at their 1995 levels. The details of the model can be found in Appendix G.1.131313In Adao, Costinot, and Donaldson (2017), the authors consider two demand systems: standard CES and “Mixed CES.” I focus on the standard CES specification. The key insight is that we can express the change in China’s welfare in period $t$, denoted by $W_{\mathrm{China},t}^{\mathrm{cf,prop}}$, as a function of all the bilateral trade flows in different periods $\left\{F_{ij,t}\right\}$ and the trade elasticity $\varepsilon$, which is estimated by $\tilde{\varepsilon}\left(\left\{F_{ij,t}\right\}\right)$. [^14] As in Section 3.1.1, I focus solely on measurement error in trade flows, implicitly assuming that all other observed data are measured without error. Hence, we can write
@@
for $t=1,...,T$ and known functions $g_{t}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\times\mathbb{R}_{++}\rightarrow\mathbb{R}$. Then, conditional on a prior distribution for the true bilateral flows $\left\{F_{ij,t}\right\}$ and a measurement error model, we can quantify uncertainty for $\left\{W_{\mathrm{China},t}^{\mathrm{cf,prop}}\right\}$.
@@ 5.1.2 Measurement Error Model and Prior
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the mirror trade data setting from Section 4.2. Since there are no zero flows in this application, the estimated posterior of interest is
@@
where $\left\{\mathring{s}_{ij}^{2}\right\}$, $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$, $\left\{\tilde{F}_{ij,t}\right\}$ and $\left\{\tilde{\mu}_{ij,t}\right\}$ are all defined in Appendix F.
@@ 5.1.3 Results
Having obtained a posterior distribution for the true trade flows given the noisy trade flows, we can now quantify uncertainty about the counterfactual predictions of interest. In Figure 5, I reproduce Figure 3 of Adao, Costinot, and Donaldson (2017), which plots the percentage change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include three different $95\%$ intervals.
The first only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, which matches the discussion in Section 3.1 and samples from the normal distribution with mean and variance equal to the GMM estimator for the trade elasticity $\varepsilon$ and its sampling variance, respectively. The resulting intervals are small for the period before the year 2000, and then slowly become wider. These are the intervals reported in Adao, Costinot, and Donaldson (2017).
The second region considers only measurement error and no estimation error in $\varepsilon$. The resulting interval is considerably wider than the interval based solely on estimation error, especially in the first few years. Finally, the third region combines estimation error and measurement error and follows Algorithm 3. The resulting bounds seem to be reasonable compositions of the bounds considering only estimation error or only measurement error. In Appendix G.3 I perform additional analyses to check the robustness of these results.
Figure 5: EB uncertainty quantification for heteroskedastic normal shocks to $\left\{\log F_{ij,t}\right\}$ for the change in China’s welfare due to the China shock. The solid blue line is the estimate as reported in Adao, Costinot, and Donaldson (2017).
5.2 Application 2: Allen and Arkolakis (2022)
@@ 5.2.1 Model and Counterfactual Question of Interest
The empirical application in Allen and Arkolakis (2022) aims to estimate the returns on investment for all highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a $1\%$ improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise only requires data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. Three of these parameters are taken from the literature and are assumed to have no uncertainty around them. The fourth, which is the congestion elasticity $\nu$, is estimated using the noisily measured traffic flow data. The details of the model can be found in Appendix H.1, but the key relation is the one that maps the average annual daily traffic (AADT) flows $\left\{F_{ij}\right\}$ to the change in welfare $W^{\mathrm{cf,prop}}$, which is
$W^{\mathrm{cf,prop}}=g\left(\left\{F_{ij}\right\},\nu\right)$
for a known function $g:\mathbb{R}_{+}^{n\left(n-1\right)}\times\mathbb{R}\rightarrow\mathbb{R}$.
@@ 5.2.2 Measurement Error Model and Prior
For this application we can again apply the default approach from Section 4. For the empirical Bayes step we can use the baseline case from Section 4.2. There are no zeros so we only have to provide an estimate of the measurement error variance $\varsigma^{2}$. Musunuru and Porter (2019) estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that Allen and Arkolakis (2022) uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$.
With $\tilde{\varsigma}^{2}=0.05$, I use Equation (8) to find a prior variance of $\tilde{s}^{2}=0.101$. This results in the following estimated posterior distribution for the true traffic flow between country $i$ and $j$, $F_{ij}$, given its noisy version $\tilde{F}_{ij}$, for $i,j=1,...,n$:
@@ 5.2.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of Allen and Arkolakis (2022) report the top ten links. For exposition, I will focus my analysis on the three best performing links. Similarly to the setting in Section 5.1, I consider scenarios with only estimation error, only measurement error, and both estimation and measurement error.
Concerning estimation error, I follow the discussion in Section 3.1 and sample from the normal distribution with mean and variance equal to the IV estimator for the congestion elasticity $\nu$ and its reported squared standard error, respectively. [^15] Here, I follow the inference method used by the authors of Allen and Arkolakis (2022) and use the clustered standard error. However, as discussed in Sanders (2025), clustering at the edge level tends to understate uncertainty relative to approaches that account for dyadic dependence in the data, and therefore yields narrower confidence intervals. Table 4 shows the $95\%$ equal-tailed intervals for the top three links. The intervals that consider just measurement error or estimation error are of similar order of magnitude, and the interval that combines them seem a sensible composition.
Point estimate Only est. error Only meas. error Est. and meas. error
Link 1 10.43 [8.33, 11.47] [8.69, 14.15] [7.86, 14.89]
Link 2 9.54 [7.02, 10.76] [7.31, 10.83] [6.60, 11.32]
Link 3 7.31 [5.05, 8.57] [6.78, 8.18] [5.30, 8.90]
Table 4: EB uncertainty quantification for the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
From a policy perspective it is of interest whether the ranking between these links can change due to estimation and measurement error. Therefore, Table 5 shows the $95\%$ intervals for the difference between link 1 and link 2, and the difference between link 2 and link 3.161616This simple exercise is intended purely for exposition. For a more formal treatment of inference on ranks, see Mogstad et al. (2024). It follows that the rankings are generally robust against estimation error and measurement error. Additional discussion and analyses can be found in Appendices H.2 and H.3.
Point estimate Only est. error Only meas. error Est. and meas. error
Link 1-Link 2 0.89 [0.61, 1.29] [0.38, 5.39] [0.38, 5.66]
Link 2-Link 3 2.23 [1.96, 2.25] [-0.05, 3.27] [0.02, 3.49]
Table 5: EB uncertainty quantification for the differences between the three links from Allen and Arkolakis (2022) with the highest return on investment. Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY).
@@ 6 Conclusion
In this paper, I provide an econometric framework to examine the effect of parameter uncertainty and measurement error for an important class of quantitative trade and spatial models. This setting departs from conventional measurement error models because the object of interest depends directly on the correctly measured data realizations, rather than on their distribution. I adopt a Bayesian approach to quantify uncertainty in counterfactual predictions, explicitly incorporating both estimation error and measurement error. I apply the framework to the settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022), and find substantial uncertainty surrounding key economic quantities in both cases. These findings highlight the importance of accounting for measurement and estimation error in counterfactual analysis.
@@
Appendix A Finding $g$ in Two Leading Classes of Models
This section informally discusses how to find the function $g$ for two leading classes of models, namely invertible models and exact hat algebra models. We are generally interested in the effect of proportional changes to the fundamentals, $X\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$. Denote these proportional changes by $X^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{X}}$. In particular, we want to find the corresponding proportional changes to the observed data, $D^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{D}}$. Our scalar prediction of interest, $\gamma$, will then be some transformation of the vector of change variables $D^{\mathrm{cf,prop}}$. That is, we consider a mapping of the form
$X^{\mathrm{cf,prop}},D,\theta,D^{\mathrm{cf,prop}}\mapsto\gamma.$
Given such a structure, it suffices to focus attention on the mapping
$X^{\mathrm{cf,prop}},D,\theta\mapsto D^{\mathrm{cf,prop}}.$ (10)
@@ A.1 Invertible Models
Redding and Rossi-Hansberg (2017) define a model to be invertible if there exists a one-to-one mapping from the observed data and structural parameter to the fundamentals. Once we have obtained the levels of the fundamentals, we can apply the proportional change of interest and find the corresponding proportional changes to the observed data. The high-level steps of this approach are:
@@
    “Back out” the levels of the fundamentals $X$ using the observed data $D$ and the structural parameter $\theta$.
@@
    Find counterfactual changes variables $D^{\mathrm{cf,prop}}$ using the counterfactual levels of the data $D\varodot D^{\mathrm{cf,prop}}$ and the baseline levels of the data $D$.
Existence of the mapping in Equation (10) follows.

### v5

@@ 6 Applications
In this section I discuss the applications in adao2017nonparametric and allen2022welfare. The application in adao2017nonparametric extends the prototypical example to a panel-data setting and recovers changes in trade costs using variation over time. The application in allen2022welfare illustrates how the proposed procedure can be extended to an economic geography framework.
6.1 Application 1: adao2017nonparametric
@@ 6.1.1 Model and Counterfactual Question of Interest
The empirical application of adao2017nonparametric investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs had stayed constant at their 1995 levels. They consider $n$ countries and $T$ time periods. The exercise I am considering is assessing the sensitivity of counterfactual predictions to measurement error in bilateral trade flows.
The counterfactual objects of interest is the change in China’s welfare, defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change where China’s trade costs are fixed at their 1995 levels. The details of the model can be found in Appendix D.1.101010In adao2017nonparametric, the authors consider two demand systems: standard CES and “Mixed CES”. I focus on the standard CES specification. The key insight is that we can express the proportional change in China’s welfare in period $t$, denoted $W_{\mathrm{China},t}^{\mathrm{cf,prop}}$, as a function of the full set of bilateral trade flows across periods, $\left\{F_{ij,t}\right\}$, and the trade elasticity, $\varepsilon$. Implementing this mapping requires multiple years of data, since answering the counterfactual question necessitates recovering the proportional change in China’s trade costs between 1995 and year $t$. This change is inferred from bilateral trade flows observed in those two years. Hence, we can write
@@
for t=1,,Tt=1,...,T and known functions $g_{t}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\times\mathbb{R}_{++}\rightarrow\mathbb{R}$. Then, conditional on a prior distribution for the true bilateral flows $\left\{F_{ij,t}\right\}$ and a measurement error model, we can quantify uncertainty for $\left\{W_{\mathrm{China},t}^{\mathrm{cf,prop}}\right\}$.
@@ 6.1.2 Measurement Error Model and Prior
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the mirror trade data setting from Section 4.2.2. Since there are no zero flows in this application, the estimated posterior of interest is
@@
where $\left\{\mathring{s}_{ij}^{2}\right\}$, $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$, $\left\{\tilde{F}_{ij,t}\right\}$ and $\left\{\tilde{\mu}_{ij,t}\right\}$ are all defined in Appendix G.
@@ 6.1.3 Results
Having obtained a posterior distribution for the true trade flows given the noisy trade flows, we can now quantify uncertainty about the counterfactual predictions of interest. In Figure 2, I reproduce Figure 3 of adao2017nonparametric, which plots the percentage change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include two $95\%$ intervals.
The first region only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, and samples from the normal distribution with mean and variance equal to the GMM estimator for the trade elasticity $\varepsilon$ and its sampling variance, respectively. The resulting intervals are small for the period before the year 2000, and then slowly become wider. These are the intervals reported in adao2017nonparametric.
The second region considers only measurement error and no estimation error in $\varepsilon$. The resulting intervals are considerably wider than the intervals based on estimation error, especially in the first few years. In Appendix D.3 I provide additional discussion and analyses.
[…]
6.2 Application 2: allen2022welfare
@@ 6.2.1 Model and Counterfactual Question of Interest
The empirical application in allen2022welfare aims to estimate the returns on investment for all highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a $1\%$ improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise only requires data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. The details of the model can be found in Appendix E.1, but the key relation is the one that maps the average annual daily traffic (AADT) flows $\left\{F_{ij}\right\}$ to the counterfactual return on investments $\left\{R_{k\ell}^{\mathrm{cf}}\right\}$, which is
$R_{k\ell}^{\mathrm{cf}}=g_{k\ell}\left(\left\{F_{ij}\right\},\theta\right)$
for known functions $g_{k\ell}:\mathbb{R}_{+}^{n\left(n-1\right)}\times\Theta\rightarrow\mathbb{R}$ for $k,\ell=1,...,n$ locations in the US Interstate Highway network.
@@ 6.2.2 Measurement Error Model and Prior
For this application we can again apply the default approach from Section 4. For the empirical Bayes step we can use the baseline case from Section 4.2.1. There are no zeros so we only have to provide an estimate of the measurement error variance $\varsigma^{2}$. musunuru2019applications estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that allen2022welfare uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$.
With $\tilde{\varsigma}^{2}=0.05$, I use Equation (9) to find a prior variance of $\tilde{s}^{2}=0.101$. This results in the following estimated posterior distribution for the true traffic flow between country $i$ and $j$, $F_{ij}$, given its noisy version $\tilde{F}_{ij}$, for i,j=1,,ni,j=1,...,n:
@@ 6.2.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of allen2022welfare report the top ten links. For exposition, I will focus my analysis on the three best performing links. Table 2 shows the $95\%$ intervals for the top three links based on Algorithm 2.
Point estimate Interval accounting for measurement error
Link 1 10.43 [8.69, 14.15]
Link 2 9.54 [7.31, 10.83]
Link 3 7.31 [6.78, 8.18]
Table 2: Uncertainty quantification for the return on investment for the three links from allen2022welfare with the highest return on investment. Returns on investment are reported as annualized decimal returns (10.43 means an annual 1043% return). Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY). The intervals based on measurement error report the 2.5th and 97.5th quantile of the estimated posterior distributions.
From a policy perspective it is of interest whether the ranking between these links can change due to measurement error. Therefore, Table 3 shows the $95\%$ intervals for the difference between link 1 and link 2, and the difference between link 2 and link 3.111111This simple exercise is intended purely for exposition. For a more formal treatment of inference on ranks, see mogstad2024inference. It follows that the rankings are generally robust against measurement error. Additional discussion and analyses can be found in Appendices E.2 and E.3.
Point estimate Interval accounting for measurement error
Link 1-Link 2 0.89 [0.38, 5.39]
Link 2-Link 3 2.23 [-0.05, 3.27]
Table 3: Uncertainty quantification for the differences in return on investment between the three links from allen2022welfare with the highest return on investment. Returns on investment are reported as annualized decimal returns (0.89 means an annual 89% return). Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY). The intervals based on measurement error report the 2.5th and 97.5th quantile of the estimated posterior distributions.
@@ 7 Conclusion
This paper develops an econometric framework for quantifying the impact of measurement error in a broad class of quantitative trade and spatial models. Unlike standard econometric models of measurement error, the counterfactual estimand in these models depends directly on the realized data rather than on the underlying distribution. I adopt an empirical Bayes approach to characterize uncertainty in counterfactual predictions and propose a default specification that can be easily implemented across a range of applications. Applying the framework to the settings in adao2017nonparametric and allen2022welfare, I find substantial uncertainty surrounding key economic outcomes. These results underscore the need to account for measurement error when using quantitative models to guide policy decisions.
@@
Appendix A Characterizing $g$ in Two Leading Classes of Models
Assumption 1 requires that, for a fixed counterfactual question and a fixed parameter value $\theta$, the counterfactual object of interest can be written as $\gamma=g\left(D,\theta\right)$, where $D$ collects the baseline inputs taken from the observed economy, and $\theta$ collects model parameters that are held fixed across baseline and counterfactual economies. This appendix discusses how to construct $g$ in two broad classes of quantitative models, namely invertible models and exact hat algebra models.
@@ A.1 Invertible Models
Following redding2017quantitative, a model is invertible if, given $\theta$, the equilibrium conditions imply a one-to-one mapping between the baseline observables $D$ and a set of latent equilibrium objects $X\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$, often called fundamentals. Formally, there exist maps $X=\mathcal{J}\left(D,\theta\right)$ and $D=\mathcal{E}\left(X,\theta\right)$, that are inverses of each other (up to standard normalizations such as a price level or numeraire). In quantitative trade and spatial models, $X$ often includes objects such as trade costs and productivity levels. For example, in the Armington model in Section 5, the fundamentals are $X=\left(\left\{\tau_{ij}\right\},\left\{\chi_{ij}\right\},\left\{\kappa_{i}\right\}\right)$.
Now consider a counterfactual experiment that specifies proportional changes to the fundamentals $X$. Denote these proportional changes by $X^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{X}}$. For example in the Armington model in Section 5, the proportional changes are $X^{\mathrm{cf,prop}}=\left(\left\{1+0.1\cdot\mathbb{I}\left\{i\neq j\right\}\right\},\left\{1\right\},\left\{1\right\}\right)$. The goal is to compute the corresponding proportional changes to the observed data, $D^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{D}}$, and then the scalar counterfactual object $\gamma$. The computational steps are:
@@
    “Invert” the model to recover the levels of the fundamentals $X$:
    $X=\mathcal{J}\left(D,\theta\right).$
@@
    Solve for counterfactual levels of the data $D\varodot D^{\mathrm{cf,prop}}$ and convert to proportional changes:
    $D\varodot D^{\mathrm{cf,prop}}=\mathcal{E}\left(X\varodot X^{\mathrm{cf,prop}},\theta\right)\quad\Rightarrow\quad D^{\mathrm{cf,prop}}=\mathcal{E}\left(X\varodot X^{\mathrm{cf,prop}},\theta\right)\oslash D,$
@@
    Compute the counterfactual object of interest:
    $\gamma=\Gamma\left(D,D^{\mathrm{cf,prop}},\theta\right).$
Composing these steps yields a function of $\left(D,\theta\right)$ alone:
$\gamma=\Gamma\left(D,\mathcal{E}\left(\mathcal{J}\left(D,\theta\right)\varodot X^{\mathrm{cf,prop}},\theta\right)\oslash D,\theta\right)\equiv g\left(D,\theta\right).$